Two angles are complementary angles if the sum of their masures is 90 degrees. Findtwo complementary angles such that one of the angles is 154 degrees lest than 3 times the oter angle.(Round two decimal place if necessary)
Answer by checkley77(12844)

(3x-154)+x=90
4x=90+154
4x=244
x=244/4
x=61 degrees.
3*61-154=183-154=29 degrees.
Proof:
61+29=90
90=90